# A Review of Bayesian Uncertainty Quantification in Deep Probabilistic Image Segmentation

### v5

@@ 8 Conclusion
Modeling the uncertainty of segmentation models is essential for accurately assessing the reliability of their predictions. Given the vast and diverse body of related literature spanning various applications and modalities, often scattered and marked by inconsistent evaluation and benchmarking practices, there is a clear need for a comprehensive and systematic overview. This work addresses that need by presenting clear definitions and a unified notation for methodologies in uncertainty modeling. Furthermore, confusion and misunderstanding around uncertainty modeling and quantification have led to diverging beliefs within the field. By structuring the methodologies by targeted uncertainty type and followed by downstream pipeline into four distinct tasks, it becomes easier for researchers to navigate the field and build upon existing work. We place particular emphasis on addressing key conceptual gaps to support both pragmatic and theoretically sound research. We also highlight the potential and broader impact of advancing uncertainty-based active learning. A noticeable trend is the adaptation of generative modeling advances to segmentation through novel, task-specific conditioning strategies for modeling ambiguity in segmentation masks. While such efforts are already underway, it remains important to closely follow progress in general generative machine learning and continue exploring its effective application to segmentation. Finally, we recommend future work to focus on improving spatial aggregation methods, standardization in benchmarking, backbone architectures, as well as the adoption of novel uncertainty estimation techniques, as many commonly used methods lack rigor. In particular, there is growing interest in distribution- and sampling-free approaches. In this manner, this review paper aims to guide researchers toward building more reliable, efficient, and uncertainty-aware segmentation models within the lightning-fast evolving field of Deep Learning-based Computer Vision.

### v8

@@ 9 Conclusion
This review synthesizes the fragmented literature on uncertainty in semantic segmentation, establishing a unified framework to bridge the gap between method developers, task specialists and applied researchers. Our analysis highlights critical challenges, including the nuanced distinction between uncertainty types, flawed assumptions in spatial aggregation, and a lack of standardization. We identify active learning, data-driven benchmarks, and transformer-based models as the most promising avenues for future research, particularly for extending novel uncertainty methods to holistic tasks like instance/panoptic segmentation. Our conclusion provides a clear method selection guide and places a strong emphasis on the need for future work to produce uncertainty that is reliable, explainable, actionable, and unbiased.